Pythagorean Theorem
Solve for the missing variable when a=10, b=? and c=14

Answer by Jstrasner(112)

hey,
So for this one we use the formula :
a^2 + b^2 = c^2
so we plug in the numbers:
10^2 + b^2 = 14^2
100 + b^2 = 196
We subtract 100 from both sides:
b^2 = 96
So b = the square root of 96
I hope this helps